OpenStudy (anonymous):

Molly has the equation 2 times 3 +6 =13 where should she put parentheses to make this equation correct?

OpenStudy (anonymous):

she cant?

OpenStudy (anonymous):

Uhh nowhere i suppose

OpenStudy (anonymous):

it always = 12 or 18

OpenStudy (anonymous):

how

OpenStudy (anonymous):

2(3+6) = 18 (2 * 3 ) + 6 = 12